# Mean-field dynamics of attractive resource interaction: From uniform to aggregated states

### v1

@@ 3.2 Analysis of the stability of limit states
Proposition 7.
For the dynamic conflict system (1) at case $n\geq 3$ with an arbitrary set of positive $c_{i}$, $i=\overline{1,n}$ and initial conditions $p_{i}^{t=0}>0$ for all $i=\overline{1,n}$ the trajectory converges to a unique fixed point $\bf p$$^{\infty}$, whose existence and analytical form are determined by Theorems 3 and 4. This equilibrium point $\bf p$$^{\infty}$ is locally asymptotically stable. Any boundary equilibrium points (where at least one coordinate $p_{i}^{\infty}=0$) are unstable.
Remark 4 (Hypothesis on Global Stability).
The unique fixed point $\bf p$$^{\infty}$, established in Proposition 7, is a global attractor for the set of all initial vectors with strictly positive coordinates.
The analytical formulas for the limit state $\bf p$$^{\infty}$ naturally lead to questions of control and optimization. The following propositions explore these criteria, which can be viewed as solutions to an optimal control problem in a deterministic mean-field system, a topic relevant to modern resource allocation challenges [26, 27].
3.3 Bifurcation analysis of limit states
The conditions derived in Theorem 4, which determine whether a component $p_{i}^{\infty}$ is zero or positive, describe a series of bifurcations (see [28, 29, 30, 31, 32, 33, 34]). The specific bifurcation occurring here is a transcritical bifurcation, where a stable interior fixed point collides with an unstable boundary fixed point and they exchange stability. This mechanism, common in ecological models [35], precisely describes how a component is eliminated from or ”switched on” in the system .
3.3.1 Codimension-one bifurcation analysis
The instability of boundary equilibria, established in the preceding analysis, is dynamically governed by bifurcations. The conditions under which a component of the steady state, $p_{i}^{\infty}$, becomes zero or positive essentially describe bifurcations. When the parameter $c_{i}$ crosses a certain threshold value, a qualitative change occurs in the system’s steady state: the coordinate $p_{i}^{\infty}$ transitions from a zero value to a positive value, or vice versa. This means that the system’s fixed point (the limiting distribution $\mathbf{p}^{\infty}$) moves from a face of the simplex (where $p_{i}=0$) into the interior of a higher-dimensional face (where $p_{i}>0$), or vice versa. Such phenomena on the boundary of the phase space are interpreted as transcritical bifurcations. Let us describe this in detail.
@@ 3.3.1 Codimension-one bifurcation analysis
Let the parameters $\{c_{2},\dots,c_{n}\}$, $n>2$ be fixed and satisfy the stability condition on the subspace:
$c_{k}>\frac{n-2}{\sum\limits_{j=2,j\neq k}^{n}\frac{1}{c_{j}}}\quad\forall k=\overline{2,n}$
@@
    If $c_{1}>c_{1}^{\text{crit}}$, the system has a unique, globally asymptotically stable attractor $\mathbf{p}^{\infty}$, in which all components are strictly positive ($p_{i}^{\infty}>0$, $\forall\ i=\overline{1,n}$).
@@
    If $c_{1}<c_{1}^{\text{crit}}$, the unique stable attractor of the system is a boundary fixed point $\tilde{\mathbf{p}}^{\infty}$, for which the first component is zero ($\tilde{p}_{1}^{\infty}=0$) while all others are positive ($\tilde{p}_{k}^{\infty}>0$ for $k\geq 2$).
@@
1. Pre-bifurcation state: $c_{1}>c_{1}^{\text{crit}}$ (Interior fixed point). Under this condition, the set of active indices at the steady state is $\mathcal{M}^{*}=\{1,\dots,n\}$. The threshold value $\Lambda$ and the first component of the vector $\mathbf{p}^{\infty}$ depend on the parameter $c_{1}$:
@@
Figure 3: Diagram of the transcritical bifurcation for the component $p_{2}^{\infty}$ as a function of the parameter $c_{2}$. This diagram visualizes the conclusions of Theorem 6 for parameters from Fig. 2.b, $c_{1}=0.8$, $c_{3}=0.9$. The vertical green dotted line at $c_{2}=0.1$ indicates the parameter value from the example in Fig. 2.b. It shows how the coordinate $p_{2}^{\infty}$ changes as the parameter $c_{2}$ is varied. For $c_{2}>c_{2}^{\text{crit}}$, an exchange of stability occurs: the trivial state $p_{2}^{\infty}=0$ loses stability, and the new state with $p_{2}^{\infty}>0$ becomes stable. This behavior is characteristic of a transcritical bifurcation.
3. Post-bifurcation state: $c_{1}<c_{1}^{\text{crit}}$ (Boundary fixed point). When $c_{1}<c_{1}^{\text{crit}}$, the condition $c_{1}>\Lambda(c_{1})$ is violated, and a fixed point with $p_{1}^{\infty}>0$ is no longer physically possible. The system evolves to a state where the set of active indices reduces to $\mathcal{M}_{\text{new}}^{*}=\{2,\dots,n\}$, with cardinality $\gamma(\mathcal{M}_{\text{new}}^{*})=n-1$. The new value $\tilde{\Lambda}$ for this subsystem on the face $p_{1}=0$ is calculated as:
@@
$\tilde{p}_{1}^{\infty}=0,\tilde{p}_{k}^{\infty}=1-\frac{\tilde{\Lambda}}{c_{k}}=1-\frac{c_{1}^{\text{crit}}}{c_{k}},k=\overline{2,n}.$
@@
which guarantees the positivity of the coordinates $\tilde{p}_{k}^{\infty}>0$. This point is the stable attractor of the system.
Thus, as the parameter $c_{1}$ passes through the critical value $c_{1}^{\text{crit}}$, a qualitative restructuring of the phase portrait occurs: the stable attractor moves from the interior of the simplex to its boundary, which corresponds to the definition of a transcritical bifurcation. ∎
3.3.2 Codimension-two bifurcation analysis
Having established the system’s dynamics for a single bifurcation parameter, we now advance the analysis to the more complex codimension-two scenario, where the qualitative behavior is considered under the variation of two control parameters, $c_{1}$ and $c_{2}$.
@@ 3.3.2 Codimension-two bifurcation analysis
Let the parameters $\{c_{3},\dots,c_{n}\}$, $n>3$, be fixed such that
$c_{k}>\frac{n-3}{\sum\limits_{j=3}^{n}\frac{1}{c_{j}}}\quad\forall k=\overline{3,n}$
@@
then $p_{2}^{\infty}=0$ and $p_{i}^{\infty}>0$ for all $i\in\{1,3,\dots,n\}$.
@@
then $p_{1}^{\infty}=p_{2}^{\infty}=0$ and $p_{i}^{\infty}>0$ for all $i\in\{3,\dots,n\}$.
Figure 4: Bifurcation diagram in the parameter space $(c_{1},c_{2})$ illustrating the conclusions of Theorem (7) for a system with $n=4$ components and fixed parameters $c_{3}=c_{4}=1$. The diagram shows the plane of control parameters partitioned into four distinct regions by the transcritical bifurcation curves $\Gamma_{1}:c_{1}\left(\frac{1}{c_{2}}+\sum_{j=3}^{n}\frac{1}{c_{j}}\right)=n-2$ and $\Gamma_{2}:c_{2}\left(\frac{1}{c_{1}}+\sum_{j=3}^{n}\frac{1}{c_{j}}\right)=n-2$.
@@
The proof is based on the result from Theorem 4, which states that for a limit fixed point $\bf p$$^{\infty}$, its $k$-th coordinate $p_{k}^{\infty}$ is positive if and only if the condition $c_{k}>\Lambda(M^{*})$ holds, where $M^{*}=\{i|p_{i}^{\infty}>0\}$ is the set of indices of active (non-zero) coordinates. The threshold value $\Lambda(M^{*})$ is defined as:
@@
1. The set of active indices is $M^{*}=\{1,2,\dots,n\}$, thus $\gamma(M^{*})=n$. The threshold value is:
@@
The condition $p_{i}^{\infty}>0$ for all $i$ requires the inequalities $c_{i}>\Lambda(M^{*})$ to hold for all $i=\overline{1,n}$. Let’s consider the condition for $c_{1}$:
@@
2. The set of active indices is $M^{*}_{1}=\{2,3,\dots,n\}$, thus $\gamma(M^{*}_{1})=n-1$. The threshold value for this subspace is:
@@
4. The set of active indices is $M^{*}_{12}=\{3,\dots,n\}$, thus $\gamma(M^{*}_{12})=n-2$. The threshold value is:
@@
The boundaries separating these parameter regions, $\Gamma_{1}:c_{1}\bigg(\frac{1}{c_{2}}+\sum\limits_{j=3}^{n}\frac{1}{c_{j}}\bigg)=n-2$ and $\Gamma_{2}:c_{2}\bigg(\frac{1}{c_{1}}+\sum\limits_{j=3}^{n}\frac{1}{c_{j}}\bigg)=n-2$, are the locations of transcritical bifurcations. The unique fixed point $\bf p$$^{\infty}$ is the locally asymptotically stable attractor of the system in each respective region.
3.3.3 Codimension-$k$ bifurcation analysis
The preceding subsections considered cases of codimension-one (one control parameter) and codimension-two (two control parameters) bifurcations. We now generalize this analysis to the case of $k$ control parameters, where $1\leq k<n$. This allows us to describe the conditions under which the system’s stable attractor collapses from a state with $n$ positive components (from the interior of the simplex) to a state on its $k$-dimensional face, where $k$ chosen components become zero.
@@ 3.3.3 Codimension-$k$ bifurcation analysis
Let the parameters $\{c_{k+1},\dots,c_{n}\}$, $n>k+1$, be fixed and satisfy the stability condition on their (nk)(n-k)-component subspace:
$c_{i}>\frac{n-k-1}{\sum\limits_{j=k+1}^{n}\frac{1}{c_{j}}}\quad\forall i=\overline{k+1,n}$
Let $c_{1},\dots,c_{k}>0$ be control (bifurcation) parameters. For any initial vector ${\bf p}^{t=0}$ with $p_{i}^{t=0}>0$ for all $i$, the trajectory ${\bf p}^{t}$ converges to a unique fixed point $\bf p$$^{\infty}$, whose structure is determined by the following cases:
If for all $i=\overline{1,k}$ the condition holds:
@@
then $p_{i}^{\infty}>0$ for all $i=\overline{1,n}$.
If for all $i=\overline{1,k}$ the condition holds:
@@
where $M^{*}_{k}=\{k+1,\dots,n\}$, then $p_{i}^{\infty}=0$ for $i=\overline{1,k}$ and $p_{j}^{\infty}>0$ for $j\in M^{*}_{k}$.
If for some non-empty subset of indices $S\subset\{1,\dots,k\}$ the conditions hold:
@@
$c_{j}>\frac{n-|S|-1}{\sum\limits_{j\in M^{*}_{S}}\frac{1}{c_{j}}}\quad\forall j\in\{1,\dots,k\}\setminus S$
where $M^{*}_{S}=\{1,\dots,n\}\setminus S$, then $p_{i}^{\infty}=0$ for $i\in S$ and $p_{j}^{\infty}>0$ for $j\in M^{*}_{S}$.
@@
The proof is a direct generalization of the proof of Theorem 7 and is based on the iterative application of the self-consistent conditions from Theorem 4. Each of the described cases corresponds to a stable attractor in the respective region of the $k$-dimensional parameter space $(c_{1},\dots,c_{k})$.
The boundaries between these regions are hypersurfaces on which transcritical bifurcations occur, analogous to the curves $\Gamma_{1},\Gamma_{2}$ depicted in Fig. 4 for the $k=2$ case. The intersection of these hypersurfaces forms a bifurcation structure of higher codimensions, which partitions the parameter phase space into regions with different numbers of active components $p_{i}^{\infty}>0$ at equilibrium. ∎
It is fundamental to note that a bifurcation of ”dimension nn” (i.e., a transition where all $n$ components vanish, $n\to 0$) is impossible within this dynamical system. The reason lies in the model’s core construction: the state vector ${\bf p}^{t}$ is a stochastic vector, which must satisfy the normalization constraint $\sum_{i=1}^{n}p_{i}^{t}=1$ for all $t\geq 0$. A limiting state ${\bf p}^{\infty}=(0,0,\dots,0)$ would imply $\sum p_{i}^{\infty}=0$, which violates this conserved quantity. Such a zero-vector is not an element of the system’s phase space (the standard (n1)(n-1)-dimensional simplex). The most extreme collapse the system can undergo is a codimension n1n-1 bifurcation, where n1n-1 components vanish simultaneously. This corresponds to the system’s attractor ”collapsing” to one of the vertices of the simplex, ${\bf p}^{\infty}={\bf e}_{k}=(0,\dots,1,\dots,0)$. This state, representing total resource aggregation into a single component, is the maximal possible aggregation and still perfectly satisfies the normalization constraint $\sum p_{i}^{\infty}=1$.
3.4 Criteria for optimizing limit coordinates
Without loss of generality, let’s assume that for all $k=\overline{1,n}$, condition 13 holds, meaning all limiting coordinates $p_{k}^{\infty}$ are strictly positive and $\gamma(M^{*})=n$. Then

### v2

@@ 3.2 Analysis of the stability of equilibria
Based on these consistent analytical considerations and extensive numerical simulations, we observe a strong pattern regarding the stability of the system for $n\geq 3$. Specifically, for the dynamic conflict system (1) with an arbitrary set of positive parameters $c_{i}$ and initial conditions $p_{i}^{t=0}>0$ for all $i=1,\ldots,n$, the trajectory consistently converges to a unique fixed point $\bf p$$^{\infty}$, whose existence and analytical form are determined by Theorems 3 and 4.
The numerical evidence compellingly suggests that this equilibrium $\bf p$$^{\infty}$ is locally asymptotically stable, while any boundary equilibria (where at least one coordinate $p_{i}^{\infty}=0$) are unstable. Furthermore, these empirical observations lead us to the hypothesis that this unique fixed point acts as a global attractor for the set of all initial vectors with strictly positive coordinates.
3.3 Bifurcation analysis of equilibria
The conditions derived in Theorem 4, which determine whether a component $p_{i}^{\infty}$ is zero or positive, describe a series of bifurcations (see [34, 35, 36, 37, 38]). It is worth noting that such structural transitions are a fundamental feature of collective dynamics in complex networks, particularly in mean–field models of coupled oscillators with mixed attractive-repulsive (conformist and contrarian) interactions [45, 46, 47, 48, 49, 50]. The specific bifurcation occurring here is a transcritical bifurcation, where a stable interior fixed point collides with an unstable boundary fixed point and they exchange stability. This mechanism, common in ecological models [39], precisely describes how a component is eliminated from or “switched on” in the system.
3.3.1 Analysis for a single control parameter
The instability of boundary equilibria, established in the preceding analysis, is dynamically governed by bifurcations. The conditions under which a component of the equilibrium, $p_{i}^{\infty}$, becomes zero or positive essentially describe bifurcations. When the parameter $c_{i}$ crosses a certain threshold value, a qualitative change occurs in the system’s asymptotic state: the coordinate $p_{i}^{\infty}$ transitions from a zero value to a positive value, or vice versa. Mathematically, this means that the system’s fixed point moves from a face of the simplex (where $p_{i}=0$) into the interior of a higher-dimensional face (where $p_{i}>0$), or vice versa. Such phenomena on the boundary of the phase space are interpreted as transcritical bifurcations.
@@ 3.3.1 Analysis for a single control parameter
Let the parameters $\{c_{2},\ldots,c_{n}\}$, $n>2$ be fixed and satisfy the stability condition on the subspace:
$c_{k}>\frac{n-2}{\sum\limits_{j=2,j\neq k}^{n}\frac{1}{c_{j}}}\quad\forall k=2,\ldots,n$
@@
    If $c_{1}>c_{1}^{\text{crit}}$, the system has a unique global attractor $\mathbf{p}^{\infty}$, in which all components are strictly positive ($p_{i}^{\infty}>0$, $\forall\ i=1,\ldots,n$).
@@
    If $c_{1}<c_{1}^{\text{crit}}$, the unique attractor of the system is a boundary fixed point $\tilde{\mathbf{p}}^{\infty}$, for which the first component is zero ($\tilde{p}_{1}^{\infty}=0$) while all others are positive ($\tilde{p}_{k}^{\infty}>0$ for $k\geq 2$).
@@
Figure 3: Diagram of the transcritical bifurcation for the component $p_{2}^{\infty}$ as a function of the parameter $c_{2}$. This diagram visualizes the conclusions of Theorem 6 for parameters from Fig. 2b, $c_{1}=0.8$, $c_{3}=0.9$. To provide a complete mathematical characterization, the diagram includes the unstable equilibrium branch in the non-physical region ($p_{2}^{\infty}<0$, red dashed curve). The vertical green dotted line at $c_{2}=0.1$ indicates the parameter value from the example in Fig. 2b. It shows how the coordinate $p_{2}^{\infty}$ changes as the parameter $c_{2}$ is varied. For $c_{2}>c_{2}^{\text{crit}}$, an exchange of stability occurs: the trivial state $p_{2}^{\infty}=0$ loses stability, and the new state with $p_{2}^{\infty}>0$ becomes stable. The intersection of these two branches and the subsequent exchange of their stability properties are definitive characteristics of a transcritical bifurcation.
@@
1. Pre-bifurcation state: $c_{1}>c_{1}^{\text{crit}}$ (Interior fixed point). Under this condition, the set of active indices at the steady state is $\mathcal{M}^{*}=\{1,\ldots,n\}$. The threshold value $\Lambda$ and the first component of the vector $\mathbf{p}^{\infty}$ depend on the parameter $c_{1}$:
@@
3. Post-bifurcation state: $c_{1}<c_{1}^{\text{crit}}$ (Boundary fixed point). When $c_{1}<c_{1}^{\text{crit}}$, the condition $c_{1}>\Lambda(c_{1})$ is violated, and a fixed point with $p_{1}^{\infty}>0$ is no longer physically possible. The system evolves to a state where the set of active indices reduces to $\mathcal{M}_{\text{new}}^{*}=\{2,\ldots,n\}$, with cardinality $\gamma(\mathcal{M}_{\text{new}}^{*})=n-1$. The new value $\tilde{\Lambda}$ for this subsystem on the face $p_{1}=0$ is calculated as:
@@
$\tilde{p}_{1}^{\infty}=0,\tilde{p}_{k}^{\infty}=1-\frac{\tilde{\Lambda}}{c_{k}}=1-\frac{c_{1}^{\text{crit}}}{c_{k}},k=2,\ldots,n.$
@@
which guarantees the positivity of the coordinates $\tilde{p}_{k}^{\infty}>0$. This point is the attractor of the system.
Thus, as the parameter $c_{1}$ passes through the critical value $c_{1}^{\text{crit}}$, a qualitative restructuring of the phase portrait occurs: the attractor moves from the interior of the simplex to its boundary, which corresponds to the definition of a transcritical bifurcation. ∎
3.3.2 Analysis for two control parameters
Having established the system’s dynamics for a single bifurcation parameter, we now advance the analysis to the scenario where two control parameters, $c_{1}$ and $c_{2}$, vary simultaneously.
@@ 3.3.2 Analysis for two control parameters
Let the parameters $\{c_{3},\ldots,c_{n}\}$, $n>3$, be fixed such that
$c_{k}>\frac{n-3}{\sum\limits_{j=3}^{n}\frac{1}{c_{j}}}\quad\forall k=3,\ldots,n$
@@
then $p_{2}^{\infty}=0$ and $p_{i}^{\infty}>0$ for all $i\in\{1,3,\ldots,n\}$.
@@
then $p_{1}^{\infty}=p_{2}^{\infty}=0$ and $p_{i}^{\infty}>0$ for all $i\in\{3,\ldots,n\}$.
Figure 4: Two-parameter bifurcation diagram in the $(c_{1},c_{2})$ plane for $n=4$ and $c_{3}=c_{4}=1$. The manifolds of transcritical bifurcations are defined by the curves $\Gamma_{1}:c_{1}\left(\frac{1}{c_{2}}+\sum_{j=3}^{n}\frac{1}{c_{j}}\right)=n-2$ and $\Gamma_{2}:c_{2}\left(\frac{1}{c_{1}}+\sum_{j=3}^{n}\frac{1}{c_{j}}\right)=n-2$. Crossing $\Gamma_{1}$ into Region II results in $p_{1}^{\infty}=0$, while crossing $\Gamma_{2}$ into Region III results in $p_{2}^{\infty}=0$. These manifolds partition the parameter space into four regions corresponding to distinct stable equilibria: Region I (interior fixed point, $p_{1}^{\infty}>0,p_{2}^{\infty}>0$); Region II (boundary fixed point, $p_{1}^{\infty}=0,p_{2}^{\infty}>0$); Region III (boundary fixed point, $p_{1}^{\infty}>0,p_{2}^{\infty}=0$); and Region IV (bounded by both curves, $p_{1}^{\infty}=p_{2}^{\infty}=0$). The curves $\Gamma_{1}$ and $\Gamma_{2}$ intersect transversally at the point $(0.5,0.5)$, where both transcritical bifurcations occur simultaneously.
@@
The proof is based on the result from Theorem 4, which states that for a fixed point $\bf p$$^{\infty}$, its $k$-th coordinate $p_{k}^{\infty}$ is positive if and only if the condition $c_{k}>\Lambda(M^{*})$ holds, where $M^{*}=\{i|p_{i}^{\infty}>0\}$ is the set of indices of active (non-zero) coordinates. The threshold value $\Lambda(M^{*})$ is defined as:
@@
1. The set of active indices is $M^{*}=\{1,2,\ldots,n\}$, thus $\gamma(M^{*})=n$. The threshold value is:
@@
The condition $p_{i}^{\infty}>0$ for all $i$ requires the inequalities $c_{i}>\Lambda(M^{*})$ to hold for all $i=1,\ldots,n$. Let’s consider the condition for $c_{1}$:
@@
2. The set of active indices is $M^{*}_{1}=\{2,3,\ldots,n\}$, thus $\gamma(M^{*}_{1})=n-1$. The threshold value for this subspace is:
@@
4. The set of active indices is $M^{*}_{12}=\{3,\ldots,n\}$, thus $\gamma(M^{*}_{12})=n-2$. The threshold value is:
@@
The boundaries separating these parameter regions, $\Gamma_{1}:c_{1}\bigg(\frac{1}{c_{2}}+\sum\limits_{j=3}^{n}\frac{1}{c_{j}}\bigg)=n-2$ and $\Gamma_{2}:c_{2}\bigg(\frac{1}{c_{1}}+\sum\limits_{j=3}^{n}\frac{1}{c_{j}}\bigg)=n-2$, are the locations of transcritical bifurcations. The unique fixed point $\bf p$$^{\infty}$ is the attractor of the system in each respective region.
3.3.3 Analysis for multiple control parameters
The preceding subsections considered cases involving one and two control parameters. We now generalize this analysis to the case of $k$ control parameters, where $1\leq k<n$. This allows us to describe the conditions under which the system’s attractor shifts from a state with $n$ positive components (from the interior of the simplex) to a state on its $k$-dimensional face, where $k$ chosen components become zero.
@@ 3.3.3 Analysis for multiple control parameters
Let the parameters $\{c_{k+1},\ldots,c_{n}\}$, $n>k+1$, be fixed and satisfy the stability condition on their (nk)(n-k)-component subspace:
$c_{i}>\frac{n-k-1}{\sum\limits_{j=k+1}^{n}\frac{1}{c_{j}}}\quad\forall i=k+1,\ldots,n.$
Let $c_{1},\ldots,c_{k}>0$ be control (bifurcation) parameters. For any initial vector ${\bf p}^{t=0}$ with $p_{i}^{t=0}>0$ for all $i$, the trajectory ${\bf p}^{t}$ converges to a unique fixed point $\bf p$$^{\infty}$, whose structure is determined by the following cases:
If for all $i=1,\ldots,k$ the condition holds:
@@
then $p_{i}^{\infty}>0$ for all $i=1,\ldots,n$.
If for all $i=1,\ldots,k$ the condition holds:
@@
where $M^{*}_{k}=\{k+1,\ldots,n\}$, then $p_{i}^{\infty}=0$ for $i=1,\ldots,k$ and $p_{j}^{\infty}>0$ for $j\in M^{*}_{k}$.
If for some non-empty subset of indices $S\subset\{1,\ldots,k\}$ the conditions hold:
@@
$c_{j}>\frac{n-|S|-1}{\sum\limits_{j\in M^{*}_{S}}\frac{1}{c_{j}}}\quad\forall j\in\{1,\ldots,k\}\setminus S$
where $M^{*}_{S}=\{1,\ldots,n\}\setminus S$, then $p_{i}^{\infty}=0$ for $i\in S$ and $p_{j}^{\infty}>0$ for $j\in M^{*}_{S}$.
@@
The proof is a direct generalization of the proof of Theorem 7 and is based on the iterative application of the self-consistent conditions from Theorem 4. Each of the described cases corresponds to an attractor in the respective region of the $k$-dimensional parameter space $(c_{1},\ldots,c_{k})$.
The boundaries between these regions are hypersurfaces on which transcritical bifurcations occur, analogous to the curves $\Gamma_{1},\Gamma_{2}$ depicted in Fig. 4. The intersection of these hypersurfaces forms a multi-parameter bifurcation structure, which partitions the parameter phase space into regions with different numbers of active components $p_{i}^{\infty}>0$ at equilibrium. ∎
It is fundamental to note that a bifurcation leading to the simultaneous vanishing of all $n$ coordinates (i.e., $p_{i}^{\infty}=0$ for all $i$) is impossible within this dynamical system. The reason lies in the model’s core mathematical construction: the state vector ${\bf p}^{t}$ is a stochastic vector, which must satisfy the normalization constraint $\sum_{i=1}^{n}p_{i}^{t}=1$ for all $t\geq 0$. The vector of equilibrium ${\bf p}^{\infty}=(0,0,\ldots,0)$ would imply $\sum_{i=1}^{n}p_{i}^{\infty}=0$, which violates this conserved quantity and does not belong to the state space (the standard (n1)(n-1)–dimensional simplex).
The most extreme topological change the system can undergo is a transition where n1n-1 coordinates vanish simultaneously. This corresponds to the system’s attractor locating at one of the vertices of the simplex, ${\bf p}^{\infty}={\bf e}_{k}=(0,\ldots,1,\ldots,0)$. In applied terms, this state represents the total aggregation of the resource into a single component, which is the maximal possible aggregation that still perfectly satisfies the normalization constraint.
3.4 Criteria for optimizing extremal coordinates
The analytical formulas for the equilibrium vector $\bf p$$^{\infty}$ naturally lead to questions of control and optimization. It is important to note that the dynamic conflict system (1) is not a gradient system; it does not explicitly minimize a potential function. However, the fact that the asymptotic state is uniquely and analytically determined by the parameters $c_{i}$ allows us to treat these parameters as control variables.
The following propositions explore these criteria, which can be viewed as solutions to an optimal control problem in a deterministic mean–field system, a topic relevant to modern resource allocation challenges [32, 33].
Without loss of generality, let’s assume that for all $k=1,\ldots,n$, condition 13 holds, meaning all coordinates $p_{k}^{\infty}$ are strictly positive and $\gamma(M^{*})=n$. Then